# WetLinks: a Large-Scale Longitudinal Starlink Dataset with Contiguous Weather Data

### 26 Feb 2024 v1

@@ II-A Starlink Overview
Customers connect to the network via their user terminal, which is colloquially known as a “Dishy”. Bidirectional communication between the terminal and the satellite makes use of $K_{u}$-Band beams. Communication between satellites and ground stations makes use of the $K_{a}$-Band. Starlink has a reconfiguration interval of 15 s, where the terminal can be rescheduled to another satellite and frequencies and routing resources can be reallocated [20]. SpaceX claims that downlink speeds up to 220 Mbps and uplinks speeds up to 25 Mbps are possible at latencies ranging from 25 to 60 ms[27]. SpaceX offers Standard and Priority service plans. Priority traffic is given network precedence over standard traffic [26] at times of high network load and congestion.
@@ IV-A Network Link Measurement
We use a three step sequential process to measure network link parameters, as shown in Fig. 3. The first two steps (the iperf3 and ping measurements) are executed every 3 minutes, the third step is executed every 6 minutes. We use iperf3 (version 3.9) [4] to measure up- and downstream throughput. The measurement of both directions is conducted in parallel using UDP with two separate measurements. While iperf3’s has a bidirectional mode, we do not use this mode since it is susceptible to CPU limitations which can lead to inaccurate measurement results. We limit the target bitrate for the uplink to 100 Mbit/s and for the downlink to 500 Mbit/s. This ensures iperf3 does not experience CPU limitations and that our measurements comply with Starlink’s fair-use policy, ensuring that the datarate is not artificially limited by Starlink. Each measurement takes 15 seconds with a five second timeout grace period.
@@ V-B Packet Loss Analysis
We next turn our attention to packet loss. Packet loss occurs frequently during our measurements. The overall statistics in Tab. II show that well over 20% of our measurements contain some packet loss. The minimum packet loss rate (PLR) was 0.4%, which corresponds to one out of 250 packets sent during the loss measurement is lost.
Fig. 5 shows the packet loss pattern over a single randomly selected 24-hour period. While the PLR is low most of the time, short spikes can be observed. These spikes typically only affect single measurements and can be interpreted as short burst losses. There is no evidence to suggest a correlation between loss events in Osnabrück and Enschede. Fig. 6 shows an ECDF of the PLR over all our measurement data split between the two sites. As the plot shows, the vast majority ($\pm$70%) of loss events concern only a single packet (PLR 0.4%). Furthermore, over 80% of measurements have a PLR $\leq$1%, and only some 3% of measurements have a PLR $\geq$5%.
@@ V-C Latency Analysis
Overall, the average bent-pipe RTT of 32-36 ms is well within the expected range from previous works [20, 12] of 30-40 ms for measurements in the EU on the 53° orbit. Our conclusion that the variance in latency is primarily caused by the bent-pipe section of the path also aligns with previous work [10, 12, 17, 19].
@@ V-D Impact of The Time of Day
Download Throughput: Fig. 8a and Fig. 8b illustrate that a typical day can be split into multiple periods. In the early morning, between 0h-5h, the median download throughput is around 230 Mbit/s in Osnabrück and 250 Mbit/s in Enschede. From 6h to approximately 15h, it decreases to around 215 Mbit/s in Osnabrück and 230 Mbit/s in Enschede. Then, a further drop can be observed until 19h, which sees the lowest throughput of 180 Mbit/s and 205 Mbit/s in Osnabrück and Enschede, respectively. Afterwards, the throughput recovers, reaching 220 Mbit/s in Osnabrück and 225 Mbit/s in Enschede around 22h.
@@
Upload Throughput: Fig. 8c and 8d indicate an increasing upload throughput between 0h-5h, then a decreasing one until 8h, and an almost constant one for the rest of the day. The minimum throughput is reached around 1h with 12.86 Mbit/s in Osnabrück, and around 3h with 13.77 Mbit/s in Enschede. The maximum throughput is reached at 5h and with 16.16 Mbit/s in Osnabrück, and at 17h with 17.66 Mbit/s in Enschede. The interquartile ranges, the whiskers and the outliers indicate a strongly fluctuating upload throughput between 0 Mbit/s and 40 Mbit/s. Interestingly, there are almost no Starlink outages at 0 Mbit/s. This implies that the Starlink uplink appears to be more stable than the downlink. Similar to the download throughput our data shows similar behaviour to previous studies [7, 12].
RTT and Packet Loss: Neither the RTT, nor the packet loss are impacted by the time of day. The median RTT is 63 ms $\pm$ 3 ms throughout the day. The median PLR is constantly 0.4%, with occasional burst losses that can occur at any time.
@@ V-E Weather Impact
Figure 10: Median download throughput rate per hour, overall both datasets Osnabrück and Enschede merged
@@
0 -10.35 12 0.08
1 -12.38 13 0.28
2 -13.73 14 4.01
3 -17.02 15 4.35
4 -20.51 16 10.34
5 -29.64 17 20.46
6 -12.37 18 24.65
7 -6.44 19 30.70
8 -5.37 20 31.72
9 -6.14 21 20.38
10 -4.59 22 3.20
11 0.27 23 -1.13
As discussed earlier in Section III, weather conditions can have an impact on satellite communication. We created a correlation matrix shown in Fig. 9 to analyze their impact on the Starlink network parameters. Rain shows a weak negative correlation of about -0.145 and -0.201 to the download throughput for the measurements in Osnabrück and Enschede respectively. All other weather factors have a correlation coefficient to the network parameters of close to zero, indicating no correlation.
@@ V-E1 The Impact of Rain
We continue our analysis with the hypothesis that rain intensity has an impact on download throughput. To verify this hypothesis, we first perform preprocessing on the data. We start by removing outliers using the interquartile range (IQR) with factor $1.5$. Next, we perform a preprocessing step to remove the time of day impact on the download throughput as shown in Fig. 8a and 8b. For that, we grouped our data into hourly buckets and performed a median correction for each bucket. Specifically, we calculated the download throughput median of all non-rainy samples of the complete dataset, and the median of the non-rainy samples within each hourly bucket. The result is shown in Fig. 10. We can clearly observe the influence of the time of day, with a difference of over 60 Mbit/s between the lowest and highest throughput. Using this data, we calculate the difference of the hourly median to the overall median and apply this difference as a correction factor ($c_{dt}$), as listed in Table III, to corresponding samples. We used a product of both datasets for this step to come up with a higher generalization grade, first by having more non-rainy samples, and second to average out the location-specific influences. A limitation of this approach is that it does not account for inter-day variance (e.g., weekday/weekend, …).
@@
Figure 12: Boxplots for different weather conditions. For each rain window 4 samples before and after the rain are used. We assume cloudy conditions before and after rain. In the no rain case, only samples with high solar radiation are used, to ensure that there are no clouds.
@@
To then verify the hypothesis of rain influence, we group the samples into rain buckets. Each bucket contains all samples with $r>u$ and $r\leq v$, where $r$ is the level of precipitation in mm, $u$ the lower threshold, and $v$ the upper threshold of the bucket. We created 7 buckets, while the last bucket contains all samples with a precipitation $>$5 mm. Additionally, we calculated a regression line, to test if an increasing rain level correlates with a change in download throughput. As the plots in Fig. 11 show, there is a clearly visible correlation, having an R value of -0.20 for Osnabrück, and an R of -0.22 for Enschede, indicating a weak correlation. A null hypothesis significance test showed a $p$-value below 0.05, which confirms our hypothesis that rain influences the download throughput significantly. The boxplots show large whiskers, indicating that factors other than rain and the time of day have a significant impact on the download throughput. These large whiskers are the main cause of the weak correlation. To reduce this noise and to further isolate the impact of rain, we added regressions through the medians of the buckets, which have an R of -0.9 and -0.98, indicating a strong linear correlation between rain and download throughput. We argue that this is a valid approach for noise reduction since the confidence intervals of the medians as well as the IQRs are small. We also tested the upload throughput against rain buckets, however, having $R$ values below 0.06 for both locations, we do not observe a significant correlation here. This may be explained by the lower bandwidth offered for upload, which provides more room for adaptation and mitigation strategies.
@@ V-E2 The Impact of Clouds
Clouds typically coincide together with rain and may also affect the download throughput. To differentiate the impact of clouds from the impact of rain, we looked at all rain periods in our dataset and extracted four samples before and after each rain window. These four samples translate to a time window of approximately 12 minutes. We assume that there is a cloudy condition shortly before and after rain periods. For comparison, we created a No Rain group, which contained all samples without rain and solar radiation higher than 300 W/m2, to ensure that there are no or only a few thin clouds. Then, we compared these four groups, as visualized in Fig. 12. When analyzing these results, we observe that the worst throughput occurs during rain with a median of 186.24 Mbit/s (95%-CI: 184.77, 187.71), which is around 14% lower than the median of the no rain condition group at 215.94 Mbit/s (95%-CI: 213.29, 218.59). Looking at the 4 sample windows before and after the rain period, we observe median throughputs of 205.16 Mbit/s before the rain (95%-CI: 203.21, 207.11) and 208.45 Mbit/s after the rain (95%-CI: 206.47, 210.43), which is almost 10 Mbit/s ($\pm$5%) lower than the median of the no rain condition group. These findings support our hypothesis that a cloudy sky also has a negative impact on the download throughput rate. However, there is a stronger negative impact of rain. To further investigate this issue, more research with an accurate labeling of the cloud situation is needed. We plan to extend our measurements in future work with this feature.
Compared to related work, our result is similar to Kassem et al.’s findings [12]. The authors observed an increase in page load time during rainy conditions. This also holds true for Ma et al. [17], who discuss the impact on upload and whose study strongly suggested that mainly the download throughput is affected by rain. A drop of around 45% in UDP download throughput for 4.1-5.2 mm rain was reported. Based on our medians for the buckets with 4-5 mm rain, we observed comparable drops of 32% and 43% for the measurements from Osnabrück and Enschede, respectively. Ma et al. [17] also indicated a correlation with temperature. Our correlation matrix, however, does not indicate such a correlation.
@@ VI Conclusion
VII Acknowledgment
We like to thank Justus Bachmann, Simon Beginn, Alexander Böckenholt, Jan Dunker, Bennet Janzen, and Malte Wehmeier for their help implementing the measurement tools and collecting the data. This work has been partially supported by the German Federal Ministry for Digital and Transport as part of the “Innovative Network Technologies” funding program (FKZ: 19OI23008C).

### 13 Mar 2024 v2

@@ II-A Starlink Overview
Customers connect to the network via their user terminal, which is colloquially known as a “Dishy”. Bidirectional communication between the terminal and the satellite makes use of $K_{u}$-Band beams. Communication between satellites and ground stations makes use of the $K_{a}$-Band. Starlink has a reconfiguration interval of 15 s, where the terminal can be rescheduled to another satellite and frequencies and routing resources can be reallocated [20]. SpaceX claims that downlink speeds up to 220 Mbit/s and uplinks speeds up to 25 Mbit/s are possible at latencies ranging from 25 to 60 ms[27]. SpaceX offers Standard and Priority service plans. Priority traffic is given network precedence over standard traffic [26] at times of high network load and congestion.
@@ IV-A Network Link Measurement
We use a three step sequential process to measure network link parameters, as shown in Fig. 3. The first two steps (the iperf3 and ping measurements) are executed approximately every 3 minutes, the third step is executed approximately every 6 minutes. We use iperf3 (version 3.9) [4] to measure up- and downstream throughput. The measurement of both directions is conducted in parallel using UDP with two separate measurements. While iperf3’s has a bidirectional mode, we do not use this mode since it is susceptible to CPU limitations which can lead to inaccurate measurement results. We limit the target bitrate for the uplink to 100 Mbit/s and for the downlink to 500 Mbit/s. This ensures iperf3 does not experience CPU limitations and that our measurements comply with Starlink’s fair-use policy, ensuring that the datarate is not artificially limited by Starlink. Each measurement takes 15 seconds with a five second timeout grace period.
@@ V-B Packet Loss Analysis
We next turn our attention to packet loss. Packet loss occurs frequently during our measurements. The overall statistics in Tab. II show that well over 25% of our measurements contain some packet loss. The minimum packet loss rate (PLR) was 0.4%, which corresponds to one out of 250 packets sent during the loss measurement is lost.
Fig. 5 shows the packet loss pattern over a single randomly selected 24-hour period. While the PLR is low most of the time, short spikes can be observed. These spikes typically only affect single measurements and can be interpreted as short burst losses. There is no evidence to suggest a correlation between loss events in Osnabrück and Enschede. Fig. 6 shows an ECDF of the PLR over all our measurement data split between the two sites. As the plot shows, the vast majority ($\pm$65%) of loss events concern only a single packet (PLR 0.4%). Furthermore, over 80% of measurements have a PLR $\leq$1%, and only some 3% of measurements have a PLR $\geq$5%.
@@ V-C Latency Analysis
Overall, the average bent-pipe RTT of 31-35 ms is well within the expected range from previous works [20, 12] of 30-40 ms for measurements in the EU on the 53° orbit. Our conclusion that the variance in latency is primarily caused by the bent-pipe section of the path also aligns with previous work [10, 12, 17, 19].
@@ V-D Impact of The Time of Day
Download Throughput: Fig. 8a and Fig. 8b illustrate that a typical day can be split into multiple periods. In the early morning, between 0h-5h, the median download throughput is around 230 Mbit/s in Osnabrück and 250 Mbit/s in Enschede. From 6h to approximately 15h, it decreases to around 210 Mbit/s in Osnabrück and 235 Mbit/s in Enschede. Then, a further drop can be observed until 19h, which sees the lowest throughput of 180 Mbit/s and 205 Mbit/s in Osnabrück and Enschede, respectively. Afterwards, the throughput recovers, reaching 210 Mbit/s in Osnabrück and 225 Mbit/s in Enschede around 22h.
@@
Upload Throughput: Fig. 8c and 8d indicate an increasing upload throughput between 0h-5h, and an almost constant one for the rest of the day. The minimum throughput is reached around 2h with 13.19 Mbit/s in Osnabrück, and 14.87 Mbit/s in Enschede. The maximum throughput is reached at 5h and with 16.23 Mbit/s in Osnabrück, and at 16h with 17.37 Mbit/s in Enschede. The interquartile ranges, the whiskers and the outliers indicate a strongly fluctuating upload throughput between 0 Mbit/s and 40 Mbit/s. Interestingly, there are almost no Starlink outages at 0 Mbit/s. This implies that the Starlink uplink appears to be more stable than the downlink. Similar to the download throughput our data shows similar behaviour to previous studies [12].
RTT and Packet Loss: Neither the RTT, nor the packet loss are impacted by the time of day. The median RTT is 62 ms $\pm$ 3 ms throughout the day. The median PLR is constantly 0.4%, with occasional burst losses that can occur at any time.
@@ V-E Weather Impact
Figure 10: Median download throughput rate per hour
@@
0 -9.24 12 0.46
1 -13.14 13 1.51
2 -12.79 14 2.26
3 -17.07 15 4.36
4 -20.55 16 10.02
5 -28.15 17 19.80
6 -14.99 18 26.23
7 -6.82 19 33.93
8 -5.27 20 34.55
9 -7.06 21 22.63
10 -2.71 22 6.93
11 -1.25 23 -0.90
As discussed earlier in Section III, weather conditions can have an impact on satellite communication. We created a correlation matrix shown in Fig. 9 to analyze their impact on the Starlink network parameters. Rain shows a weak negative correlation of about -0.165 and -0.192 to the download throughput for the measurements in Osnabrück and Enschede, respectively. All other weather factors have a correlation coefficient to the network parameters of close to zero, indicating no correlation.
@@ V-E1 The Impact of Rain
We continue our analysis with the hypothesis that rain intensity has an impact on download throughput. To verify this hypothesis, we first perform preprocessing on the data. We start by removing outliers using the interquartile range (IQR) with factor $1.5$. Next, we perform a preprocessing step to remove the time of day impact on the download throughput, as shown in Fig. 8a and 8b. For that, we grouped our data into hourly buckets and performed a median correction for each bucket. Specifically, we calculated the download throughput median of all non-rainy samples of the complete dataset, and the median of the non-rainy samples within each hourly bucket. The result is shown in Fig. 10. We can clearly observe the influence of the time of day, which is consistent at both measurement sites (same timezone), with a difference of over 60 Mbit/s between the lowest and highest throughput. Using this data, we calculate the difference of the hourly median to the overall median and apply this difference as a correction factor ($c_{dt}$), as listed in Table III, to corresponding samples. We used a product of both datasets for this step to come up with a more robust correction factor, by having more non-rainy samples. A limitation of this approach is that it does not account for inter-day variance (e.g., weekday/weekend, …).
@@
Figure 12: Boxplots for different weather conditions. For each rain window, 4 samples before and after the rain are used. We assume cloudy conditions before and after rain. In the no rain case, only samples with high solar radiation are used, to ensure that there are no clouds.
@@
To then verify the hypothesis of rain influence, we group the samples into rain buckets. Each bucket contains all samples with $r>u$ and $r\leq v$, where $r$ is the level of precipitation in mm, $u$ the lower threshold, and $v$ the upper threshold of the bucket. We created 7 buckets, while the last bucket contains all samples with a precipitation $>$5 mm. Additionally, we calculated a regression line, to test if an increasing rain level correlates with a change in download throughput. As the plots in Fig. 11 show, there is a clearly visible correlation, having an R value of -0.21 at both sites, indicating a weak correlation. A null hypothesis significance test showed a $p$-value below 0.05, which confirms our hypothesis that rain influences the download throughput significantly. The boxplots show large whiskers, indicating that factors other than rain and the time of day have a significant impact on the download throughput. These large whiskers are the main cause of the weak correlation. To reduce this noise and to further isolate the impact of rain, we added regressions through the medians of the buckets, which have an R of -0.84 and -0.98, indicating a strong linear correlation between rain and download throughput. We argue that this is a valid approach for noise reduction since the confidence intervals of the medians as well as the IQRs are small. We also tested the upload throughput against rain buckets, however, having $R$ values below 0.05 for both locations, we do not observe a significant correlation here. This may be explained by the lower bandwidth offered for upload, which provides more room for adaptation and mitigation strategies.
@@ V-E2 The Impact of Clouds
Clouds typically coincide together with rain and may also affect the download throughput. To differentiate the impact of clouds from the impact of rain, we looked at all rain periods in our dataset and extracted four samples before and after each rain window. These four samples translate to a time window of approximately 12 minutes. We assume that there is a cloudy condition shortly before and after rain periods. For comparison, we created a No Rain group, which contained all samples without rain and solar radiation higher than 300 W/m2, to ensure that there are no or only a few thin clouds. Then, we compared these four groups, as visualized in Fig. 12. When analyzing these results, we observe that the worst throughput occurs during rain with a median of 181.52 Mbit/s (95%-CI: 180.09, 182.96), which is around 17% lower than the median of the no rain condition group at 218.12 Mbit/s (95%-CI: 215.60, 220.64). Looking at the 4 sample windows before and after the rain period, we observe median throughputs of 208.08 Mbit/s before the rain (95%-CI: 206.36, 209.80) and 208.47 Mbit/s after the rain (95%-CI: 206.73, 210.21), which is almost 10 Mbit/s lower than the median of the no rain condition group. These findings support our hypothesis that a cloudy sky also has a negative impact on the download throughput rate. However, there is a stronger negative impact of rain. To further investigate this issue, more research with an accurate labeling of the cloud situation is needed. We plan to extend our measurements in future work with this feature.
Compared to related work, our result is similar to Kassem et al.’s findings [12]. The authors observed an increase in page load time during rainy conditions. This also holds true for Ma et al. [17], who discuss the impact on upload and whose study strongly suggested that mainly the download throughput is affected by rain. A drop of around 45% in UDP download throughput for 4.1-5.2 mm rain was reported. Based on our medians for the buckets with 4-5 mm rain, we observed comparable drops of 31% and 30% for the measurements from Osnabrück and Enschede, respectively. Ma et al. [17] also indicated a correlation with temperature. Our correlation matrix, however, does not indicate such a correlation.